graph the following by plotting points
y=x+6

Answer by ewatrrr(24785)

 
Hi,
graph the equation by plotting points
y=x+6
y = 0 +6
0 = x +6
Pt(0,6) and Pt(-6,0) on this line
Plotting points and connecting with a Line